# Online Multivariate Regularized Distributional Regression for High-dimensional Probabilistic Electricity Price Forecasting

### v1

@@ 2.1 Preliminaries and Setting
Online Coordinate Descent for Regularized Linear Regression
@@
$\hat{\beta}_{j}\xleftarrow{}\frac{S\left(\bm{\mathrm{H}}\left[j\right]-\bm{\mathrm{G}}\left[j,:\right]\bm{\mathrm{\beta}}+\bm{\mathrm{G}}\left[j,j\right]\hat{\beta}_{j},\lambda\right)}{\bm{\mathrm{G}}\left[j,j\right]}.$ (1)
where $S(x,\lambda)=\operatorname{sign}(x)\max(x-\lambda)$ is the so-called soft-threshold function. Coordinate descent is commonly solved on a decreasing grid of regularization strengths $\bm{\mathrm{\lambda}}$ on an exponential grid from $\lambda_{\text{max}}=\max{}\lvert\bm{\mathrm{G}}_{n+1}\rvert$. Algorithm 1 presents the full fitting process. A more detailed treatment of online coordinate descent can be found in Messner and Pinson (2019); Hirsch et al. (2024).
@@ 8                   Update $\hat{\beta}_{j,\lambda}$ according to Equation 1
Output: $\widehat{\bm{\mathrm{\beta}}}_{n+1}=\left(\hat{\beta}_{j,\lambda},...\right)^{\top}$ for all $\lambda\in\bm{\mathrm{\lambda}}$
@@
Univariate Distributional Regression Setting
Distributional regression aims to model the conditional distribution parameters of the univariate response vector $\bm{\mathrm{y}}=(y_{1},...,y_{N})\in\mathbb{R}^{N\times 1}$, conditional on the covariate or explanatory data in $\bm{\mathrm{X}}\in\mathbb{R}^{N\times J}$ by adopting a parametric distribution $\bm{\mathrm{y}}\sim\mathcal{F}(\Theta)$, where $\mathbf{\Theta}=(\bm{\mathrm{\theta}}_{1},...,\bm{\mathrm{\theta}}_{K})$ is a tuple of $K$ distribution parameters $\bm{\mathrm{\theta}}_{k}=(\theta_{k1},...,\theta_{kN})$. Each of the distribution parameters are linked to the covariate data through a known, twice differentiable link function $g_{k}(\cdot)$, leading to:
Multivariate Distributional Regression Setting
Let us appreciate here that this formulation is rather general. In practice, the different distribution parameters $\bm{\mathrm{\theta}}_{1},...,\bm{\mathrm{\theta}}_{K}$ can have many different shapes. Take, e.g. the multivariate $t$-distribution, parameterized using the Cholesky factor of the precision matrix $\bm{\mathrm{\Sigma}}=\bm{\mathrm{L}}^{\top}\bm{\mathrm{L}}$, denoted as $t_{D}(\bm{\mathrm{\theta}}_{1},\bm{\mathrm{\theta}}_{2},\bm{\mathrm{\theta}}_{3})\Leftrightarrow t_{D}(\bm{\mathrm{\mu}},\bm{\mathrm{L}},\bm{\mathrm{\nu}})$. Then $\bm{\mathrm{\mu}}$ is a ${N\times D}$ matrix, $\bm{\mathrm{L}}$ is a ${N\times D\times D}$ cube (of which each vertical slice is a triangular matrix) and $\bm{\mathrm{\nu}}$ is a $N\times 1$ vector. Accordingly, the index set $\mathcal{M}_{k}$ of coordinates spans $\mathcal{M}_{1}=\{1,..,D\}$, $\mathcal{M}_{2}=\{(1,1),...,(D,D)\}$ and $\mathcal{M}_{3}=\{1\}$ and its cardinality is given by the product of the parameter’s dimensions beyond $N$. The general setting introduced here includes the Gaussian multivariate distributional regression introduced by Muschinski et al. (2022), the Copula-based multivariate distributional by Kock and Klein (2023) and the MCD-based additive covariance models by Gioia et al. (2022). The general estimation principle of repeatedly iterating through the distribution parameters until convergence translates in the multivariate case. However, we now introduce an additional inner cycle through all coordinates of the currently active distribution parameter. The exact estimation algorithm will be introduced in Section 2.3 and the following Section 2.2 briefly discusses different options to parameterize the covariance respectively precision matrix.
2.2 Parameterization of the Precision Matrix
Covariance Matrix Modeling Strategies
Covariance and precision matrices are commonly modeled in three distinct fashions: through the use of sparse or graphical estimators (e.g. the graphical LASSO, see Friedman et al., 2008), through covariance functions (see e.g. Browell et al., 2022) or the matrices’ elements arc modeled conditional on explanatory variables, usually through an unconstrained parametrization (see e.g. Pourahmadi, 2011; Muschinski et al., 2022; Salinas et al., 2019). Our approach falls into the third category. To save computational costs, we parameterize the distributions in terms of the inverse covariance matrix $\bm{\mathrm{\Sigma}}^{-1}=\bm{\mathrm{\Omega}}$. This allows to avoid matrix inversion in the evaluation of the (log-) likelihood function. To ensure the positive definiteness of the scale matrix, we propose two unconstrained parameterizations of the precision matrix.
Cholesky-Decomposition
The CD has been introduced as suitable covariance parameterization by Pourahmadi (2011) in the context of GLMs. For the distributional regression framework, Muschinski et al. (2022); Kock and Klein (2023) employ the (modified) CD for the covariance matrix. For
$\bm{\mathrm{\Sigma}}=\bm{\mathrm{A}}\bm{\mathrm{A}}^{\top}$ $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$ (5)
Muschinski et al. (2022) parametrize the normal distribution in terms of $A^{-1}$ and Kock and Klein (2023) choose $A$. Additionally the modified Cholesky-decomposition (MCD) can be used, see e.g. Pourahmadi (2011); Muschinski et al. (2022); Gioia et al. (2022). For the CD to yield a positive semi-definite matrix, we require the diagonal of $\bm{\mathrm{A}}$ to be positive, which can be enforced by employing a log-link function. The lower diagonal of $\bm{\mathrm{A}}$ is unconstrained.
Low-Rank Approximation
The low-rank approximation (LRA) for the precision matrix has been proposed by Salinas et al. (2019) and März (2022) in the context of high-dimensional Gaussian processes and distributional gradient boosted trees. The LRA is defined as
where $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{D})$ and $\bm{\mathrm{V}}$ is a $D\times R$ matrix of rank $R$. The advantage of the LRA is that the dimensions of the parameters $\bm{\mathrm{A}}$ and $\bm{\mathrm{V}}$ scale linearly with the dimension $D$, however, the partial derivatives of the multivariate Gaussian and $t$-distribution with respect to the coordinates of $\bm{\mathrm{A}}$ and $\bm{\mathrm{V}}$ require inversion of the precision matrix. To ensure positive-definiteness for the LRA, we require the non-zero elements of $\bm{\mathrm{A}}$ to be positive, while $\bm{\mathrm{V}}$ is unconstrained. These requirements can easily be satisfied by choosing the log-link function for $\bm{\mathrm{A}}$.
Overview
Rigby and Stasinopoulos (2005) introduce iteratively reweighted least squares for generalized additive models for location, shape and scale (GAMLSS). The RS algorithm consists of two nested loops, in which we cycle repeatedly through the distribution parameters and run a weighted fit of the score vector $u$ on the design matrix $\bm{\mathrm{X}}$ using the diagonal weight matrix $\bm{\mathrm{W}}$. The following paragraphs introduce the scoring vector and weights, the algorithm and the necessary modifications to move from a univariate case to the multivariate case.
Scoring and Weights
The score vector is defined as
$\bm{\mathrm{u}}=\frac{\partial\ell}{\partial\eta}$ (7)
$\bm{\mathrm{z}}=g\big{(}\hat{\bm{\mathrm{\theta}}}\big{)}+\frac{\partial\ell}{\partial\eta}\bm{\mathrm{W}}^{-1}\Leftrightarrow\bm{\mathrm{z}}=\bm{\mathrm{\eta}}+\frac{\partial\ell}{\partial\eta}\bm{\mathrm{W}}^{-1}$ (8)
where the weights are defined as:
for Newton-Raphson and Fisher’s scoring respectively. In the GLM, Fisher’s scoring and Newton-Raphson scoring coincide for the canonical link functions in the exponential family. However, for the scale and shape parameters, this is not necessarily the case anymore (for a detailed treatment of GLMs and estimation theory, see e.g. Lange et al., 2010). In the original GAMLSS, Rigby and Stasinopoulos (2005) use Fisher’s scoring. Our approach generally uses Newton-Raphson scoring for the multivariate case, since the derivation of the expected value of second derivatives can be intractable, especially for more complex parameterizations of the precision matrix.
Mix-and-Match Newton-Raphson Scoring
Newton-Raphson scoring requires the partial derivatives of the log-likelihood function with respect to the predictors. While many previous works on distributional regression employ Newton-Raphson scoring, each derive the partial derivatives for specific combinations of distribution function and link function only (see e.g. O’Neill and Burke, 2023; Muschinski et al., 2022). To facilitate the computational implementation in an mix-and-match fashion, we propose to use the first and second derivative of the log-likelihood with respect to the parameter and the first and second derivative of the link function and relate both to the necessary derivatives for Newton-Raphson scoring using the equalities given in the following Lemma 2.1, which allow for the simple utilization of arbitrary link functions and efficient calculation of working vector and weight matrices.
Param. Dim. Param. Dim. Param. Dim.
Multivariate Gaussian $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$ $N\times\text{triangular}(D\times D)$
Multivariate Gaussian $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=\bm{\mathrm{A}}+\bm{\mathrm{V}}\bm{\mathrm{V}}^{\top}$ $N\times\text{diag}(D)$, $D\times r$
Lemma 2.1
The proof is straight-forward and utilizes the chain and quotient rules and can be found in Appendix A.3.
We provide the necessary first and second partial derivatives of the log-likelihood with respect to the distribution parameter’s coordinates, ${\partial\ell}/{\partial\theta}$ and ${\partial^{2}\ell}/{\partial\theta^{2}}$, for all parameters for the multivariate normal and multivariate $t$-distribution given in Table 1. The derivation can be found in Appendix A.4 and Appendix A.5.
Link functions
Commonly, the log-link is used for scale parameters in the univariate and multivariate distributional regression models. However, the inverse transformation through the exponential is prone to yield extreme values (see e.g. Ziel, 2022). Narajewski and Ziel (2020) propose the use of the so-called logident link function, defined as
$g(x)=\operatorname{LogIdent}(x)=\begin{cases}\log(x)&\text{if $x<1$}\\ x-1&\text{else.}\end{cases}$ (12)
However, the function is not continuously twice differentiable. It can be made twice differentiable by a sigmoid spline on the non-differentiable part and then defined as
$g(x)=\operatorname{DifferentiableLogIdent}(x)=\begin{cases}\log(x)&\text{if $x<1$}\\ (1-f(x))\log(x)+f(x)(x-1)&\text{else.}\end{cases}$ (13)
where $f(x)=1/\exp(-k(x-1))$ is a sigmoid-function that ensures a smooth transition at $x=1$ for some constant $k$. Alternatively one can use the inverse softplus function as link function. The softplus and its inverse function are popular activation functions for neural networks (see e.g. Dubey et al., 2022; Dugas et al., 2000) and have been used in GLMs by Wiemann et al. (2024). The softplus and its inverse are defined as
$g(x)=\operatorname{InverseSoftPlus}(x)$ $=\log(\exp(x)-1)$ (14)
$g^{-1}(x)=\operatorname{SoftPlus}(x)$ $=\log(1+\exp(x))$ (15)
the inverse softplus function has (almost) linear behavior for large values of $x$ and log-like behavior for small $x$. Figure 3 compares the link functions
Figure 3: Comparison of the Log, Sqrt, LogIdent and InverseSoftPlus link and inverse functions. All map a distribution parameter to the positive real line $(0,\infty)$. The graceful, almost linear behavior of the inverse transformation of the InverseSoftPlus is clearly visible compared to the square root and log link.

### v2

@@ 2 Online Multivariate Distribution Regression
2.1 Distributional Regression Setting
Starting from the univariate case, distributional regression aims to model the conditional distribution parameters of the response vector $\bm{\mathrm{y}}=(y_{1},...,y_{N})\in\mathbb{R}^{N\times 1}$, conditional on the covariate or explanatory data in $\bm{\mathrm{X}}\in\mathbb{R}^{N\times J}$ by adopting a parametric distribution $\bm{\mathrm{y}}\sim\mathcal{F}(\Theta)$, where $\mathbf{\Theta}=(\bm{\mathrm{\theta}}_{1},...,\bm{\mathrm{\theta}}_{K})$ is a tuple of $K$ distribution parameters $\bm{\mathrm{\theta}}_{k}=(\theta_{k1},...,\theta_{kN})$. Each of the distribution parameters are linked to the covariate data through a known, twice differentiable link function $g_{k}(\cdot)$, leading to:
This formulation is rather general. We note two important points here:
  • 1.
    The distribution parameters are subject to constraints. This especially holds for the scale (respectively precision) matrix, which is often required to be positive semi-definite. To ensure this holds, unconstrained parameterizations such as the (modified) Cholesky-decomposition or the low-rank approximation are often used (Pourahmadi, 2011; Muschinski et al., 2022; Salinas et al., 2019). The parameterization of the scale matrix will be discussed in Section 2.2.
  • 2.
    In practice, the different distribution parameters $\bm{\mathrm{\theta}}_{1},...,\bm{\mathrm{\theta}}_{K}$ can have many different shapes. Take, e.g. the multivariate $t$-distribution, parameterized using the Cholesky factor of the precision matrix $\bm{\mathrm{\Sigma}}=\bm{\mathrm{L}}^{\top}\bm{\mathrm{L}}$, denoted as $t_{D}(\bm{\mathrm{\theta}}_{1},\bm{\mathrm{\theta}}_{2},\bm{\mathrm{\theta}}_{3})\Leftrightarrow t_{D}(\bm{\mathrm{\mu}},\bm{\mathrm{L}},\bm{\mathrm{\nu}})$. Then $\bm{\mathrm{\mu}}$ is a $N\times D$} matrix, $\bm{\mathrm{L}}$ is a $N\times D\times D$} cube (of which each vertical slice is a triangular matrix) and $\bm{\mathrm{\nu}}$ is a $N\times 1$ vector. Accordingly, the index set $\mathcal{M}_{k}$ of coordinates spans $\mathcal{M}_{1}=\{1,..,D\}$, $\mathcal{M}_{2}=\{(1,1),...,(D,D)\}$ and $\mathcal{M}_{3}=\{1\}$ and its cardinality is given by the product of the parameter’s dimension beyond $N$.
The general setting introduced here includes the Gaussian multivariate distributional regression introduced by Muschinski et al. (2022), the Copula-based multivariate distributional by Kock and Klein (2023) and the MCD-based additive covariance models by Gioia et al. (2022). The general estimation principle of repeatedly iterating through the distribution parameters until convergence translates to the multivariate case. However, in this setting we introduce an additional inner cycle through all coordinates of the currently active distribution parameter. The exact estimation algorithm will be introduced in Section 2.3 and the following Section 2.2 briefly discusses different options to parameterize the covariance respectively precision matrix.
2.2 Parameterization of the Precision Matrix
To ensure the positive definiteness of the scale matrix, we propose three unconstrained parameterizations. To save computational costs, we parameterize the distributions in terms of the inverse covariance matrix $\bm{\mathrm{\Sigma}}^{-1}=\bm{\mathrm{\Omega}}$. This allows to avoid matrix inversion in the evaluation of the (log-) likelihood function. We briefly review first the (modified) Cholesky decompositions (CD resp. MCD), which have been proposed in the context of multivariate distributional regression by e.g. Pourahmadi (2011); Muschinski et al. (2022); Kock and Klein (2023). Additionally, we employ a low-rank approximation of the precision matrix, which has been used by Salinas et al. (2019) in the context of high-dimensional Gaussian processes and by März (2022) and O’Malley et al. (2023) in the context of distributional gradient boosted random forests.
The Cholesky-decomposition (CD) of the covariance matrix $\bm{\mathrm{\Sigma}}$ and the precision matrix $\bm{\mathrm{\Omega}}$ are defined as:
$\bm{\mathrm{\Sigma}}=\bm{\mathrm{A}}\bm{\mathrm{A}}^{\top}$ and $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$ (4)
Muschinski et al. (2022) parametrize the normal distribution in terms of $\bm{\mathrm{A}}^{-1}$ and Kock and Klein (2023) choose $\bm{\mathrm{A}}$. Additionally the modified Cholesky-decomposition (MCD) can be used:
$\bm{\mathrm{\Sigma}}=(\bm{\mathrm{L}}^{-1})^{\top}\bm{\mathrm{D}}\bm{\mathrm{L}}^{-1}$ and $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}$ (5)
For the CD to yield a positive definite matrix, we require the diagonal of $\bm{\mathrm{A}}$ to be positive, which can be enforced by employing a log-link function. The lower diagonal of $\bm{\mathrm{A}}$ is unconstrained. The same holds for the MCD, where $\bm{\mathrm{D}}$ is a diagonal matrix with positive entries and $\bm{\mathrm{L}}$ is a unit lower triangular matrix with ones on the diagonal. Note that $\bm{\mathrm{A}}^{-1}=\bm{\mathrm{D}}^{-1/2}\bm{\mathrm{L}}$. The low-rank approximation (LRA) is defined as
where $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{D})$ and $\bm{\mathrm{V}}$ is a $D\times R$ matrix of rank $R$. The advantage of the LRA is that the dimensions of the parameters $\bm{\mathrm{A}}$ and $\bm{\mathrm{V}}$ scale linearly with the dimension $D$, however, the partial derivatives of the multivariate Gaussian and $t$-distribution with respect to the coordinates of $\bm{\mathrm{A}}$ and $\bm{\mathrm{V}}$ require inversion of the precision matrix. To ensure positive-definiteness for the LRA, we require the non-zero elements of $\bm{\mathrm{A}}$ to be positive, while $\bm{\mathrm{V}}$ is unconstrained. These requirements can easily be satisfied by choosing the log-link or the square root link function for $\bm{\mathrm{A}}$.
Rigby and Stasinopoulos (2005) introduce iteratively reweighted least squares for the estimation of GAMLSS models. The RS algorithm consists of two nested loops, in which we cycle repeatedly through the distribution parameters and run a weighted fit of the score vector $\bm{\mathrm{z}}$ on the design matrix $\bm{\mathrm{X}}$ using the diagonal weight matrix $\bm{\mathrm{W}}$. The following paragraphs introduce the scoring vector and weights, the algorithm and the necessary modifications to move from a univariate case to the multivariate case.
The score vector is defined as
$\bm{\mathrm{u}}=\frac{\partial\ell}{\partial\eta},$ (7)
$\bm{\mathrm{z}}=g\big(\hat{\bm{\mathrm{\theta}}}\big)+\frac{\partial\ell}{\partial\eta}\bm{\mathrm{W}}^{-1}\Leftrightarrow\bm{\mathrm{z}}=\bm{\mathrm{\eta}}+\frac{\partial\ell}{\partial\eta}\bm{\mathrm{W}}^{-1},$ (8)
where the weights are defined as:
for Newton-Raphson and Fisher’s scoring respectively. Algorithm 1 shows the high-level algorithm for fitting distributional regression models using IRLS. To move from the univariate to the multivariate case, we need to cycle through all coordinates $m\in\mathcal{M}_{k}$ of the currently active distribution parameter $\bm{\mathrm{\theta}}_{k}$ in the inner loop and calculate the score vector $\bm{\mathrm{u}}_{km}$, working vector $\bm{\mathrm{z}}_{km}$ and weight matrix $\bm{\mathrm{W}}_{km}$ for each coordinate and run the weighted regression on $\bm{\mathrm{z}}_{km}$ on $\bm{\mathrm{X}}$. To allow for high-dimensional covariate spaces for each coordinate of the distribution parameter, we employ online coordinate descent and LASSO penalties for the weighted regression step, as introduced in Section 2.4. Section 2.5 will give a detailed view on the online version, also including the necessary details on multivariate distributions, model selection, and online updates.
1for outer iterations $i=0,$... until convergence do
2   Iterate through all distribution parameters.
3 forall $k\in\mathcal{K}$ do
4      Fit one distribution parameter.
5    for inner iterations $r=0,1,$... until convergence do
6         Evaluate three steps:
7
  1. 1.
    Evaluate $\bm{\mathrm{u}}_{k}$, $\bm{\mathrm{W}}_{k}$ and $\bm{\mathrm{z}}_{k}$ using Equations (7), (8) and (9).
  2. 2.
    Weighted regression of $\bm{\mathrm{z}}_{k}$ on $\bm{\mathrm{X}}_{k}$ using $\bm{\mathrm{W}}_{k}$ and yield $\hat{\bm{\mathrm{\beta}}}_{k}$
  3. 3.
    Evaluate convergence and end if converged.
8        -0.5em
9
10  Evaluate convergence and end if converged.
Output: $\widehat{\bm{\mathrm{\beta}}}_{k}$ and $\widehat{\bm{\mathrm{\Theta}}}=(\widehat{\bm{\mathrm{\theta}}}_{0},...,\widehat{\bm{\mathrm{\theta}}}_{p})$
Algorithm 1 High-level description of the IRLS algorithm for estimating distributional regression models (Rigby and Stasinopoulos, 2005; Stasinopoulos et al., 2024).
In the GLM, Fisher’s scoring and Newton-Raphson scoring coincide for the canonical link functions in the exponential family. However, for the scale and shape parameters, this is not necessarily the case anymore (for a detailed treatment of GLMs and estimation theory, see e.g. Lange et al., 2010). In the original GAMLSS, Rigby and Stasinopoulos (2005) use Fisher’s scoring. Our approach generally uses Newton-Raphson scoring for the multivariate case, since the derivation of the expected value of second derivatives can be intractable, especially for more complex parameterizations of the precision matrix. Newton-Raphson scoring requires the partial derivatives of the log-likelihood function with respect to the predictors. While many previous works on distributional regression employ Newton-Raphson scoring, each derive the partial derivatives for specific combinations of distribution function and link function only (see e.g. O’Neill and Burke, 2023; Muschinski et al., 2022). To facilitate the computational implementation in an mix-and-match fashion, we propose to use the first and second derivative of the log-likelihood with respect to the parameter and the first and second derivative of the link function and relate both to the necessary derivatives for Newton-Raphson scoring using the equalities given in the following Lemma 2.1, which allow for the simple utilization of arbitrary link functions and efficient calculation of working vector and weight matrices.
Lemma 2.1
The proof is straight-forward and utilizes the chain and quotient rules and can be found in A.2. We provide the necessary first and second partial derivatives of the log-likelihood with respect to the distribution parameter’s coordinates, ${\partial\ell}/{\partial\theta}$ and ${\partial^{2}\ell}/{\partial\theta^{2}}$, for all parameters for the multivariate normal and multivariate $t$-distribution given in Table 2. The derivation can be found in A.3 and A.4.
Param. Dim. Param. Dim. Param. Dim.
Multivariate Gaussian $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$ $N\times\text{triangular}(D\times D)$ - -
Multivariate Gaussian $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}$ $N\times\text{triangular}(D\times D)$, $N\times\text{diag}(D)$ - -
Multivariate Gaussian $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=\bm{\mathrm{A}}+\bm{\mathrm{V}}\bm{\mathrm{V}}^{\top}$ $N\times\text{diag}(D)$, $D\times r$ - -
Multivariate-$t$ $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}$ $N\times\text{triangular}(D\times D)$, $N\times\text{diag}(D)$ $\nu$ $N\times 1$
2.4 Online Coordinate Descent for LASSO
@@
$\hat{\beta}_{j}\xleftarrow{}\frac{S\left(\bm{\mathrm{h}}\left[j\right]-\bm{\mathrm{G}}\left[j,:\right]\bm{\mathrm{\beta}}+\bm{\mathrm{G}}\left[j,j\right]\hat{\beta}_{j},\lambda\right)}{\bm{\mathrm{G}}\left[j,j\right]}$ (12)
where $S(x,\lambda)=\operatorname{sign}(x)\max(x-\lambda)$ is the so-called soft-threshold function. Coordinate descent is commonly solved on a decreasing grid of regularization strengths $\bm{\mathrm{\lambda}}$ on an exponential grid from $\lambda_{\text{max}}=\max{}\lvert\bm{\mathrm{G}}_{n+1}\rvert$. Algorithm 2 presents the full fitting process. The formulation can be easily extended to solve ElasticNet and Ridge Regression problems, however, in this work we restrict ourselves to the LASSO estimation, as the sparse solution allows for easy use of information criteria (IC) for model selection. The implementation in the ondil Python package supports the ElasticNet as well as various extensions such as box-constrained coefficients and early stopping. A more detailed treatment of online coordinate descent can be found in Messner and Pinson (2019) and Hirsch et al. (2024).
@@ 8         Update $\hat{\beta}_{j,\lambda}$ according to Equation 12
[…]